a. What is the distribution of X? X ~ N(,)

b. Find the probability that a randomly selected person's IQ is over 107.  Round your answer to 4 decimal places.

c. A school offers special services for all children in the bottom 3% for IQ scores. What is the highest IQ score a child can have and still receive special services? (Note: Use the closest value in your z table). Round your answer to 1 decimal place.

d. Find the Inter Quartile Range (IQR) for IQ scores. Round your answers to whole numbers.
Q1: 
Q3: 
IQR: 
Answer #1

ms. Mean (M) = 120 std devo) = 15 107-120 5 a) XNN (120, 15) b) p (x > 107) = P(x-m 107-120 ) Plz>-0.87) = P125-0.87) = 1-0.1Inter quartile vange (PQR) = Oz-8 = 150-110 - (20